OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°. a. (3squareroot2, 315°), (-3squareroot2 , 135°) b. (3squareroot2, 225°), (-3squareroot2 , 45°) c. (3squareroot2, 45°), (-3squareroot2 , 225°) d. (3squareroot2 , 135°), (-3squareroot2 , 315°)

OpenStudy (anonymous):

someone!

OpenStudy (anonymous):

\[ \left\{3 \sqrt{2} \cos (315 {}^{\circ}),3 \sqrt{2} \sin (315 {}^{\circ})\right\}=\{3,-3\}\\ \left\{-3 \sqrt{2} \cos (135 {}^{\circ}),-3 \sqrt{2} \sin (135 {}^{\circ})\right\}=\{3,-3\} \] So the answer is a